# REPARO: Compositional 3D Assets Generation with Differentiable 3D Layout Alignment

### v1

@@ Abstract
Traditional image-to-3D models often struggle with scenes containing multiple objects due to biases and occlusion complexities. To address this challenge, we present REPARO, a novel approach for compositional 3D asset generation from single images. REPARO employs a two-step process: first, it extracts individual objects from the scene and reconstructs their 3D meshes using off-the-shelf image-to-3D models; then, it optimizes the layout of these meshes through differentiable rendering techniques, ensuring coherent scene composition. By integrating optimal transport-based long-range appearance loss term and high-level semantic loss term in the differentiable rendering, REPARO can effectively recover the layout of 3D assets. The proposed method can significantly enhance object independence, detail accuracy, and overall scene coherence. Extensive evaluation of multi-object scenes demonstrates that our REPARO offers a comprehensive approach to address the complexities of multi-object 3D scene generation from single images. The demo page have been available at https://reparo2024.github.io/.
Figure 1: Compositional 3D assets generated by our REPARO.
@@ 1 Introduction
The swift advancement of the 2D diffusion models and technologies, like neural rendering [34] and Score Distillation Sampling (SDS)  [39], have greatly enhanced the efficient generation of high-quality 3D assets. This progress empowers creators in augmented reality (AR) / virtual reality (VR), gaming, filmmaking, and more [47, 24, 20]. Advanced simulation systems used in real-world engineering and entertainment scenarios often require settings that involve the combination of multiple objects. However, traditional text-to-3D [39, 25, 58] and image-to-3D [33, 49, 30, 27] generation methods, while effective in creating single-object scenes, face significant challenges in reconstruct complex multi-object scenes accurately from limited references. This is because most 3D samples feature a single centrally located object in the dataset, and image-to-3D models typically recenter the object during preprocessing, resulting in an inherent center bias in these models. Thus, they often struggle with the textures and details of occluded objects, which can be mistakenly perceived as a single entity, leading to issues that separate but occluded assets "sticking" together, as shown in Figure 2 (d).
This challenge prompted us to consider: how to build a 3D generation pipeline that can independently manage multi-object assets? To overcome the above limitations, our approach involves decomposing the scene into individual assets initially and then reassembling and globally optimizing their layout. Such a two-step process significantly enhances the fidelity and independence of each object within the overall scene, addressing the inherent difficulties in handling complex occlusions and interactions among multiple objects.
Another key difficulty of multi-object compositional 3D generation is understanding and implementing context and relational dynamics between objects within a scene. This limitation stems not only from the intricacy of the geometric representation of 3D content [34, 16, 52] but also the fact that existing 3D datasets focus predominantly on a single object with very limited examples of multiple interacting objects. Recently, some studies have begun to explore the multi-object compositional 3D generation. Recent work [12] synthesizes compositional 3D scenes from either scene graphs or unstructured text descriptions. However, given the limited information conveyed by text descriptions, we take single-image as guidance and integrate differentiable 3D layout alignment.
During the assembly and layout alignment phase, our method integrates each object into a unified coordinate system. We employ differentiable rendering techniques to refine the spatial arrangement of each object. Notably, instead of using the typical pixel-wise $L_{2}$ loss common in differentiable rendering, we utilize the Optimal Transport (OT) algorithm to facilitate a global one-to-one pixel matching between the rendered image and the reference image. Moreover, we propose incorporating high-level semantic features into the loss function of differentiable rendering, enhancing the model’s ability to maintain context and coherence among the objects.
Additionally, the ability to create interactive environments where users can manipulate individual elements without disrupting the overall scene’s structure and realism is also vital. Traditional 3D generation models often produce assets tied to a single mesh, complicating the task for game designers and graphic artists who need to segment these meshes using specialized tools—a process that does not ensure the quality of the segmented assets. Our targeted approach facilitates the individual generation of each asset, which can then be assembled as per the original design. This method eliminates the need for complex segmentation and joining operations, as our meshes are generated at the instance level within the scene, thus simplifying the workflow and enhancing usability for end-users.
To summarize, our methodology REPARO comprises two distinct phases: individual object reconstruction and combined scene optimization. Initially, we isolate each target object from the provided image and complete any occluded sections, then reconstruct their 3D meshes employing off-and-shelf 3D generation models. Subsequently, we refine the spatial configuration of these meshes through differentiable rendering, aiming to achieve a coherent layout of the scene. REPARO enhances object independence and detail accuracy, especially in occluded sections. Meanwhile, REPARO maintains not only the visual-spatial arrangement but also contextual consistency across the reconstructed scene.
@@
To assess our approach, we selected 20 scenes from the GSO dataset [7], each containing multiple objects (ranging from 2 to 6 sub-objects) for quantitative testing. Additionally, we gathered 20 more images featuring multiple objects for qualitative analysis. This comprehensive evaluation demonstrates that our method, REPARO, offers significant improvements over previous techniques in ensuring the quality of assets, managing multiple assets, and processing occlusion.
Our contributions are summarized as follows:
@@
    We introduce REPARO, a novel approach that decomposes complex scenes into individual 3D components and then reconstructs them into accurate, high-quality multi-object scenes, enhancing object independence and detail.
@@
    We propose a long-range appearance loss term for the differentiable rendering to optimize the layout of 3D assets. This loss term uses the optimal transport algorithm to find a pixel match between the rendered image and the reference image, thus obtaining long-range correspondences. Then, we utilize RGB color, depth value, and position distance at the same time.
@@
    We propose a high-level semantic loss term for the layout alignment, significantly improving the contextual coherence and interactive potential of generated 3D scenes.
@@ 2 Related Work
Text-to-3D generation. Early exploration in 3D generation based on text [62, 34, 35, 55, 36, 41, 15] relied on a large-scale pre-trained model of text and image, CLIP [41], to optimize representation as the prior. In light of the promising capability of diffusion models, some works [39, 25, 1, 58, 42, 21, 56, 43] introduce Score Distillation Sampling (SDS) to narrow the gap between novel view images rendered from 3D representation and diffusion prior. Recent works [48, 65, 3, 23] have used 3D Gaussian Splatting [16] instead of NeRF [34] for 3D generation, due to the fact that this efficient 3D representation shows excellent quality in reconstruction tasks.
Image-to-3D generation. Differing from the ambiguity and diversity associated with textual description, 3D generation based on a single image predominantly concentrates on maintaining geometric and textural consistency between the generated assets and the input images. Some early works [33, 49] used 2D diffusion models as prior knowledge to guide the training of 3D representation of the object in input images. However, with the enrichment of 3D data [6, 5], a series of adapting fine-tuned diffusion models [29, 44, 28, 40, 27, 60, 67] represented by Zero-1-to-3 [30] have been continuously developed. Subsequently, further advancements [31, 45, 46, 32] have been proposed, which offer improved quality and multi-view consistency. More recently, a new paradigm [14, 63] getting rid of diffusion models relying on camera view as condition has emerged, mainly based on transformer architecture [53] and triplane [34] representation.
Compositional 3D generation.
In contrast to single-object reconstruction, generating complex 3D assets that contain multiple objects requires meticulous attention to detail not only for individual objects but also for their spatial relationships and occlusions. Responding to the challenges in achieving realistic and coherent compositions, related studies are mostly split into two sub-tasks constructing local single objects and modeling global scenes.
To configure the layout of the scene and place objects, the 3D geometric surface model [11] is constructed through compositing the retrieved and interactive segmented 3D meshes. Part-aware generative model [50] is proposed for editable 3D shape synthesis. Bounding boxes [38] and text prompts [57] are utilized for scene component creation. Explicit Gaussian radiance field [54] is used as guidance of object compositions. Scene graphs [12] are used to generate compositional 3D scenes by exploiting node and edge information. Large language models (LLMs) are integrated into pipeline [26] as a 3D layout interpreter. Method [2] applied spatially-aware score distillation sampling to guide the positions of 3D assets, while work [9] learned 3D NeRF layouts for text-prompted scenes. Recent work [64] employed a 2D semantic layout map as an input condition to generate compositional 3D scenes. Meanwhile model [66] leverages Gaussian Splatting and LLMs to create complex 3D scenes from textual descriptions. Differently, our method focuses on directly generating multiple high-quality 3D assets from a single image containing complex compositional objects, while the positions of assets are aligned with the layout of the input image.
@@ 3 Method
Figure 2: Qualitative comparison of generated 3D assets of the single object and multiple objects. (b) and (c) are generated 3D assets of the single object assets using DreamGaussian [48]; (d) is a 3D asset of multiple objects generated by DreamGaussian [48]; and (e) is our result.
The field of 3D generation is advancing rapidly, with numerous image-to-3D models [48, 51] now capable of generating 3D assets from a single image. However, these models typically focus on individual objects, making it challenging to apply them to scenes containing multiple objects. To address this limitation, we propose REPARO for compositional 3D object generation from a single image. Our approach consists of two steps: (1) Extracting each target object from the given image and reconstructing their 3D meshes using off-the-shelf 3D reconstruction models. (2) Optimizing the parameters of each mesh through differentiable rendering so as to align their layout. In the following section, we provide a detailed explanation of these two steps.
@@ 3.1 Reconstruction of Each Single Object
Single Object Reconstruction. After that, we employ off-the-shelf image-to-3D models, such as Dreamgaussian [48], to generate the corresponding 3D assets for each obtained object.
@@ 3.2.1 Long-range Appearance Loss Term
Optimal transport theory describes the following problem: suppose there are $N$ suppliers and $M$ demanders within a region. The $i$-th supplier holds $s_{i}$ units of goods, and the $j$-th demander needs $d_{j}$ units of goods. The transportation cost per unit of goods from $i$-th supplier to $j$-th demander is $c_{ij}$. The goal of the optimal transport algorithm is to find a transportation matrix $T=\{T_{ij}\mid i=1,\ldots,N;j=1,\ldots,M\}$ that minimizes the total transportation cost:
$\min_{T}\sum_{i=1}^{N}\sum_{j=1}^{M}T_{ij}c_{ij},\quad s.t.\quad\sum_{i=1}^{N}{T_{ij}}=d_{j},\quad\sum_{j=1}^{M}{T_{ij}}=s_{i},\quad\sum_{i=1}^{N}{s_{i}}=\sum_{i=j}^{M}{d_{i}},\quad T_{ij}>0.$ (3)
@@
$L_{a}(I,I^{ref})=\frac{1}{N}\sum_{i}^{N}\left(\alpha\cdot\|I_{i}-I^{ref}_{\sigma(i)}\|_{2}+\beta\cdot\|D_{i}-D^{ref}_{\sigma(i)}\|_{2}+\gamma\cdot\|p_{i}-p_{\sigma(i)}\|_{2}\right).$ (5)
@@ 3.2.2 High-level Semantic Loss Term
where $\lambda$ is a hyper-parameter to adjust the weight between appearance and semantic loss term. In practice, we align the layout by optimizing the object’s translation parameter $t=\{x,y,z\}$ and rotation parameter $r$. Thanks to the proposed loss function, we can get the layout of the compositional 3D assets through differentiable rendering.
@@ 4 Experiments
4.1 Implementation details
In the reconstruction of 3D assets from a single object, we utilize SAM-ViT-H [18] as the segmentation model to obtain masks corresponding to each bounding box. In the preprocessing module, we extract images of the objects using these masks and subsequently resize and crop the images to center the objects. We then apply the Stable Diffusion-based inpainting model  [^1] https://github.com/lkwq007/stablediffusion-infinity to address occlusions in the objects. During the generation of 3D assets from a single object, we employ the DreamGaussian [48] and TripoSR [51]. For layout alignment, we use Nvdiffrast [19] as our differentiable rendering framework. During the differentiable rendering process, we optimize the translation $x,y,z$, and rotation $r$ parameters of each 3D asset using gradient descent. Specifically, we employ the Adam optimizer [17] with a learning rate of $0.02$ and a weight decay of $0.999$. The total number of iterations is set to $500$. Throughout the experiment, both the reference and rendered images are maintained at a resolution of $256\times 256$. In the long-range appearance loss function $L_{a}$, we estimate the depth of the images using DPT-DINOv2-base [37]. For pixel matching, we adopt Sinkhorn divergences [4] as an efficient approximation of the optimal transport algorithm and use a GPU implementation provided by GemoLoss [10] with the parameter $\epsilon$ set to 0.01. In the high-level semantic loss function $L_{s}$, we extract image features using DINOv2-base [37] and compute the loss using features of the last hidden state. $L_{s}$ is subjected to a warm-up period of 200 iterations. In Eq. 9, we set $\lambda$ to 0.9. All experiments were conducted on RTX 3090 GPUs with the alignment process for each sample taking approximately 5 minutes and utilizing 4GB of memory.
4.2 Main results
Table 1: Performance of different 3D generation models for compositional 3D assets generation.
Method CLIP score $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@ 4.2 Main results
DreamGaussian-based REPARO (ours) 0.833 17.391 0.827 0.232
TripoSR-based REPARO (ours) 0.823 17.778 0.865 0.216
Figure 4: Qualitative comparison with different image-to-3D generation models.
Quantitative experiment. Table 1 compares the performance (CLIP score, PSNR, SSIM, and LPIPS) of various methods for reconstructing 3D assets of multiple objects using one reference image. DreamGaussian-based REPARO and Triposr-based REPARO, our proposed methods, show significant improvements over baseline methods (DreamGaussian [48], Wonder3D [32], and LRM [14]). DreamGaussian-based REPARO achieves 83.3% CLIP score, 17.391 PSNR, 0.827 SSIM, and 0.232 LPIPS, indicating better alignment between reconstructed 3D assets and ground-truth 3D assets. Triposr-based REPARO exhibits superior performance in PSNR (17.778) and SSIM (0.865) and the lowest LPIPS (0.216), highlighting its exceptional quality in image reconstruction and perceptual similarity. These results show that better reconstruction quality on single objects is beneficial for structural similarity and perceptual similarity during differentiable rendering.

### v2

@@ Abstract
footnotetext: Equal contribution. Corresponding authors.
Traditional image-to-3D models often struggle with scenes containing multiple objects due to biases and occlusion complexities. To address this challenge, we present REPARO, a novel approach for compositional 3D asset generation from single images. REPARO employs a two-step process: first, it extracts individual objects from the scene and reconstructs their 3D meshes using image-to-3D models; then, it optimizes the layout of these meshes through differentiable rendering techniques, ensuring coherent scene composition. By integrating optimal transport-based long-range appearance loss term and high-level semantic loss term in the differentiable rendering, REPARO can effectively recover the layout of 3D assets. The proposed method can significantly enhance object independence, detail accuracy, and overall scene coherence. Extensive evaluation of multi-object scenes demonstrates that our REPARO offers a comprehensive approach to address the complexities of multi-object 3D scene generation from single images. The demo have been available at https://reparo-3d.github.io/.
Figure 1: Qualitative comparison of generated 3D assets of the single object and multiple objects. (b) and (c) are generated 3D assets of the single object assets using DreamGaussian [51]; (d) is a 3D asset of multiple objects generated by DreamGaussian [51]; (e) is our result.
@@ 1 Introduction
The field of 3D content creation has advanced rapidly due to progress in reconstruction methods [35, 16] and generative frameworks [40]. Automated 3D content generation enables creators in augmented reality (AR), virtual reality (VR), gaming, and filmmaking [50, 25, 20] by significantly reducing manual labor requirements.
Recent research in 3D generation primarily focuses on optimization-based 2D approaches, particularly text-to-3D [40, 26, 60] and image-to-3D [34, 52, 31, 27] models. For example, Dreamfusion [40] overcomes 3D data scarcity by distilling geometry and appearance from 2D diffusion models; and DreamGaussian [51] integrates 3D Gaussian Splatting into generative pipelines with mesh extraction and texture refinement. However, despite their effectiveness in single-object scenes, these methods face challenges in reconstructing multi-object scenes. This limitation stems from dataset biases: most 3D training samples feature centrally aligned single objects, and preprocessing pipelines often re-center inputs, thereby introducing inherent positional bias. Consequently, occluded objects are frequently misrepresented as fused entities, leading to erroneously merged assets in rendered outputs, as shown in Figure 1 (d). Furthermore, typical 3D generation models output assets as monolithic mesh representations. This forces users to rely on error-prone post-processing steps to segment individual object mesh, which does not ensure the quality of the segmented assets.
To address this challenge, we propose REPARO, a compositional 3D generation pipeline capable of generating multiple objects from a single input image. Our approach decomposes the scene into discrete assets, then systematically reassembles and globally optimizes their spatial arrangement. This methodology not only leverages the strengths of off-the-shelf models (the ability to generate single objects with high fidelity) but also addresses inherent challenges in the multi-object scene with complex occlusions and interactions.
Specifically, our pipeline comprises two core stages. First, individual objects are isolated from the input image through cropping and inpainting to generate context-complete single-object images, which are processed using off-the-shelf image-to-3D models [41, 33, 54] to produce high-fidelity 3D assets. Second, the 3D assets are assembled into a cohesive scene and geometrically aligned with the input image. After placing all objects within a unified coordinate system, we employ differentiable rendering to optimize the spatial arrangement of each asset. During rendering, the pose parameters (translation, rotation, and scale) of each 3D asset are treated as optimizable variables, which are iteratively refined via gradient descent using a loss function. Unlike conventional pixel-wise loss function [3], we introduce an optimal transport (OT)-based long-range appearance loss that integrates RGB color, depth, and positional information to constrain global structural relationships between objects. Therefore, this loss can match the rendered image and the reference image. Additionally, a high-level semantic loss is introduced to ensure semantic alignment and coherence between the synthesized scene and the original image context using high-level visual features.
Through these two stages, individual-object reconstruction and combined-scene optimization, our REPARO enhances object independence and geometric detail accuracy, particularly for occluded regions. Simultaneously, it preserves both spatial arrangement fidelity and semantic-geometric consistency across the synthesized scene. Critically, our method obviates the need for post hoc mesh segmentation or manual joining operations by generating instance-level meshes directly within the scene context, streamlining the 3D asset creation workflow and improving end-user usability.
@@
To evaluate our method, we conducted experiments on 20 multi-object scenes from the Google Scanned Objects (GSO) dataset [9] (containing 2–6 objects per scene) and 20 additional in-the-wild images for qualitative analysis. Quantitative and qualitative comparisons demonstrate that REPARO offers significant improvements over previous techniques in ensuring the quality of assets, managing multiple assets, and processing occlusion. Our contributions are summarized as follows:
@@
    A two-stage compositional generation Pipeline: We propose REPARO, a framework to decouple multi-object 3D generation into two stages: (1) per-object high-fidelity reconstruction via inpainting-aware cropping and off-the-shelf models, and (2) global scene optimization using novel differentiable rendering. This structure ensures object-level geometric independence while preserving scene-level contextual integrity.
@@
    Optimal transport-driven layout loss: We design a long-range appearance loss based on optimal transport (OT) theory, which enforces multi-modal constraints (RGB, depth, positional channels) to align 3D asset layouts with the reference image. By solving for global pixel correspondences, OT mitigates local minima issues inherent in the pixel-wise loss.
@@
    Semantic loss: We further introduce a semantic alignment loss that leverages high-level visual features to refine object placements. This loss ensures synthesized scenes adhere to the original image’s semantic context.
@@ 2 Related work
Text-to-3D generation. Early exploration in 3D generation based on text [63, 35, 36, 56, 37, 44, 15] relied on a large-scale pre-trained model of text and image, CLIP [44], to optimize representation as the prior. In light of the promising capability of diffusion models, some works [40, 26, 1, 60, 45, 21, 57, 46] introduce Score Distillation Sampling (SDS) to narrow the gap between novel view images rendered from 3D representation and diffusion prior. Recent works [51, 65, 4, 24] have used 3D Gaussian Splatting [16] instead of NeRF [35] for 3D generation, due to the fact that this efficient 3D representation shows excellent quality [66, 43, 23] in reconstruction tasks.
Image-to-3D generation. Differing from the ambiguity and diversity associated with textual description, 3D generation based on a single image predominantly concentrates on maintaining geometric and textural consistency between the generated assets and the input images. Some early works [34, 52] used 2D diffusion models as prior knowledge to guide the training of 3D representation of the object in input images. However, with the enrichment of 3D data [6, 7], a series of adapting fine-tuned diffusion models [30, 47, 29, 42, 27, 61, 68] represented by Zero-1-to-3 [31] have been continuously developed. Subsequently, further advancements [32, 49, 48, 33] have been proposed, which offer improved quality and multi-view consistency.
Compositional 3D generation. Unlike single-object reconstruction, multi-object 3D generation requires attention to both local details and global spatial relationships. Existing approaches address this through two main paradigms: (1) Single-object construction using geometric surface models [10] and editable generative models [53]; (2) Scene composition employing bounding boxes [39], text prompts [59], Gaussian radiance fields [55], and scene graphs [11]. Emerging approaches integrate LLMs for layout interpretation [28], leverage 2D semantic maps [64], or combine Gaussian Splatting with LLMs [67]. Distinct from text/layout-based methods, our approach directly generates multiple high-quality 3D assets from a single compositional image while preserving input-aligned spatial arrangements.
@@ 3 Method
The field of 3D generation is advancing rapidly, with numerous image-to-3D models [51, 54] capable of generating 3D assets from a single image. However, these models typically focus on individual objects, making it challenging to apply them to scenes containing multiple objects. To address this limitation, we propose REPARO for compositional 3D object generation from a single image. Our approach consists of two steps: (1) Extracting each target object from the given image and reconstructing their 3D meshes using off-the-shelf 3D reconstruction models. (2) Optimizing the parameters of each mesh through differentiable rendering to align their layout. In the following section, we provide a detailed explanation of these two steps.
@@ 3.1 Reconstruction of Each Single Object
Single object reconstruction. After that, we employ off-the-shelf image-to-3D models, such as Dreamgaussian [51] and TripoSR [54], to generate the corresponding 3D assets for each obtained object.
@@ 3.2.1 Long-range Appearance Loss Term
Optimal transport theory describes the following problem: suppose there are $N$ suppliers and $M$ demanders within a region. The $i$-th supplier holds $s_{i}$ units of goods, and the $j$-th demander needs $d_{j}$ units of goods. The transportation cost per unit of goods from $i$-th supplier to $j$-th demander is $c_{ij}$. The goal of the optimal transport algorithm is to find a transportation matrix $T=\{T_{ij}>0\mid i=1,\ldots,N;j=1,\ldots,M\}$ that minimizes the total transportation cost:
$\min_{T}$ $\sum_{i=1}^{N}\sum_{j=1}^{M}T_{ij}c_{ij},\quad T_{ij}>0,$ (3)
s.t. $\sum_{i=1}^{N}{T_{ij}}=d_{j},\quad\sum_{j=1}^{M}{T_{ij}}=s_{i},\quad\sum_{i=1}^{N}{s_{i}}=\sum_{j=1}^{M}{d_{j}}.$
@@
$L_{a}(I,I^{ref})=\frac{1}{N}\sum_{i}^{N}c_{i\sigma(i)}.$ (5)
@@ 3.2.2 High-level Semantic Loss Term
where $\lambda$ is a hyper-parameter to adjust the weight between appearance and semantic loss term. In practice, we align the layout by optimizing the object’s translation parameter $t=\{\text{trans\_x},\text{trans\_y},\text{trans\_z}\}$ and scale parameter $s=\{\text{scale\_x},\text{scale\_y},\text{scale\_z}\}$. We exclude the rotation parameter $r$ of assets from the optimization process, as assets generated via image-to-3D models are loaded into the scene with the same orientation as in the input image, eliminating the need for rotation optimization. Thanks to the proposed loss function, we can get the layout of the compositional 3D assets through differentiable rendering.
@@ 4 Experiments
Figure 3: GPU memory usage and approximate elapsed time for optimization of compositional assets layout under different parameter settings.
[…]
Method CLIP $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@
[…]
Table 2: Runtime and memory usage of each stage.
Stage VRAM Time
SAM Segmentation 6 GB <1s<1s
Inpainting 8 GB 20 $s$
[…]
Layout Alignment 6 GB 90 $s$
[…]
Table 3: Comparison with representative methods on perceptual quality and resource usage.
Method CLIP $\uparrow$ GPU Time
DreamGaussian 0.807 8 GB 120 $s$
TripoSR 0.795 6 GB <1s<1s
Wonder3D 0.801 16 GB 180 $s$
LRM 0.812 14 GB 20 $s$
[…]
4.1 Implementation details
In the reconstruction of 3D assets from a single object, we utilize SAM-ViT-H [18] as the segmentation model to obtain masks corresponding to each bounding box. In the preprocessing module, we extract images of the objects using these masks and subsequently resize and crop the images to center the objects. We then apply the Stable Diffusion-based inpainting model [14] to address occlusions in the objects. During the generation of 3D assets from a single object, we employ the DreamGaussian [51] and TripoSR [54]. For layout alignment, we use Nvdiffrast [19] as our differentiable rendering framework. During the differentiable rendering process, we optimize the translation $x,y,z$, and rotation $r$ parameters of each 3D asset using gradient descent. Specifically, we employ the Adam optimizer [17] with a learning rate of $0.02$ and a weight decay of $0.999$. The total number of iterations is set to $500$. All experiments were conducted on RTX 3090 GPUs, and as shown in Figure 3, optimizing parameter $t$ alone takes approximately 20 seconds and utilizes 4GB of memory, while optimizing both parameters $t$ and $s$ to their optimal states requires about 90 seconds and utilizes 6GB of memory. We report per-stage runtime and memory in Table 3. All modules run on $\leq$8GB VRAM, covering SAM, inpainting, 3D generation, and layout alignment. Table 3 compares our method with baselines in runtime and memory. REPARO achieves better performance under similar or lower resource budgets.
Figure 4: Qualitative comparison with different image-to-3D generation models. Given an input image, previous methods produce inaccurate textures and geometry with noticeable artifacts. Our method generates high-quality, high-fidelity compositional assets with the correct spatial layout.
Table 4: Comparison of different loss functions. (a) DreamGaussian-based REPARO. $L_{a}$ $L_{s}$ CLIP $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ 0.833 17.296 0.826 0.233 0.809 17.820 0.849 0.210 0.833 17.279 0.826 0.234 (b) TripoSR-based REPARO. $L_{a}$ $L_{s}$ CLIP $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ 0.822 17.765 0.865 0.216 0.813 17.906 0.867 0.212 0.822 17.751 0.865 0.216
Table 5: Comparison of optimization parameters. (a) DreamGaussian-based REPARO. T R S CLIP $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ 0.832 17.369 0.826 0.232 0.826 16.906 0.820 0.243 0.833 17.279 0.826 0.234 0.831 17.122 0.823 0.241 (b) TripoSR-based REPARO. T R S CLIP $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ 0.819 17.797 0.865 0.216 0.822 17.768 0.864 0.219 0.822 17.751 0.865 0.216 0.826 17.752 0.864 0.219
Table 6: Comparison of information in appearance loss $L_{a}$. (a) DreamGaussian-based REPARO. Info in $L_{a}$ OT CLIP $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RGB $\times$ 0.800 18.054 0.853 0.206 RGB 0.828 17.009 0.821 0.245 RGBXY 0.828 17.069 0.821 0.243 RGBDXY 0.833 17.279 0.826 0.234 (b) TripoSR-based REPARO. Info in $L_{a}$ OT CLIP $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RGB $\times$ 0.774 17.918 0.868 0.216 RGB 0.819 17.837 0.865 0.215 RGBXY 0.823 17.847 0.865 0.213 RGBDXY 0.822 17.751 0.865 0.216
4.2 Main results
Quantitative experiment. Table 7 compares the performance (CLIP score, PSNR, SSIM, and LPIPS) of various methods for reconstructing 3D assets of multiple objects using one reference image. DreamGaussian-based REPARO and Triposr-based REPARO, our proposed methods, show significant improvements over baseline methods (DreamGaussian [51], Wonder3D [33], and LRM [13]). DreamGaussian-based REPARO achieves 83.3% CLIP score, 17.279 PSNR, 0.826 SSIM, and 0.234 LPIPS, indicating better alignment between reconstructed 3D assets and ground-truth 3D assets. Triposr-based REPARO exhibits superior performance in PSNR (17.751) and SSIM (0.865) and the lowest LPIPS (0.216), highlighting its exceptional quality in image reconstruction and perceptual similarity. These results show that better reconstruction quality on single objects is beneficial for structural similarity and perceptual similarity during differentiable rendering.